OpenStudy (anonymous):

The length of a rectangle is twice the width. The area is 338 yd^. Find the length and width.

OpenStudy (anonymous):

1. Area equation 2. equation relating base and width. you got it

OpenStudy (anonymous):

I am looking for the answer

OpenStudy (agreene):

338 = 2x + x 338 = 3x \[\frac{338}{3}=x\] \[w = \frac{338}{3}\] \[l = \frac{676}{3}\]

OpenStudy (anonymous):

so w=112.666 and l = 225.333, correct?

OpenStudy (agreene):

If you're an engineer, I suppose.

OpenStudy (anonymous):

this is for ALG. II, it is a pain in the a$$

OpenStudy (anonymous):

AGL

OpenStudy (agreene):

Well, the fractions are more correct, because they are more exacting.

OpenStudy (agreene):

but, most maths teachers tell you to round, because most math teachers are idiots.

OpenStudy (anonymous):

that is so very true. so w=112.0, and l=225.0, lol

OpenStudy (agreene):

:) if you need to round, yes

OpenStudy (anonymous):

I have never been very good at all with word problems, they always screw me up.